Oranges are $1.30 a pound , apples are $.90 a pound and a mixture of them costs $1.10 a pound. If 15 lbs I of oranges are in the mix how many pounds of apples are needed?
Answer by lwsshak3(11628)

Oranges are $1.30 a pound , apples are $.90 a pound and a mixture of them costs $1.10 a pound. If 15 lbs I of oranges are in the mix how many pounds of apples are needed?
let x=lbs of apples needed for the mixture
15+x=lbs of apples needed for the mixture
15+x=weight of mixture
1.30*15+.90x=1.10(15+x)
19.50+.90x=16.50+1.10x
.20x=3
x=15
lbs of apples needed for the mixture=15